OpenStudy (anonymous):

Find a polynomial P(x) having real coefficients, with the degree and zeroes indicated. (Assume the lead coefficient is 1.)" roots of 1-(sqrt3), 1+(sqrt3), and 7-i

OpenStudy (calculusfunctions):

We have 2 distinct real roots of\[1-\sqrt{3}\]\[1+\sqrt{3}\]Also we have 2 conjugate complex roots 7 + i and 7 - i. Remember that complex (imaginary) roots always occur in conjugate pairs. Thus we have a quartic (4th degree) polynomial due to the fact that we have a total of 4 roots (2 real and 2 imaginary).

OpenStudy (anonymous):

Ok how do I set up to get the equation

OpenStudy (calculusfunctions):

Given the roots, the corresponding factors are\[(x -1+ \sqrt{3}),(x -1-\sqrt{3}),(x -7+i),and(x -7-i)\]. Now use these factors to write the equation in standard form.

OpenStudy (calculusfunctions):

@acestro504 ?? Do you understand?

OpenStudy (anonymous):

Sorry for the delay. I got it. Thanks very much.